OpenStudy (anonymous):

1. What is the area of a circle with a diameter of 8? (1 point) 4pi symbol 8pi symbol 16pi symbol 64pi symbol

OpenStudy (anonymous):

area of a circle is given by \[=\pi r^{2}\]

OpenStudy (anonymous):

The equation is pir^2 and r=.5d

OpenStudy (anonymous):

where r is radius

OpenStudy (anonymous):

So you could also rewrite it as "pi"(.5d)^2

OpenStudy (anonymous):

so first convert the diameter into radius and put it into that formula

OpenStudy (anonymous):

^ :D

OpenStudy (lgbasallote):

it's always funny to watch two users race with two different solutions. usually confuses askers...but still funny

OpenStudy (anonymous):

lol^(:

OpenStudy (anonymous):

Ahaha sorry I wasn't really looking while I was typing xD Good race though ;D

OpenStudy (anonymous):

aha itz ok thanx everyone!(:

OpenStudy (anonymous):

sooooo what is the answer lol???

OpenStudy (anonymous):

r=d/2 r=8/2=4 so area=pi(4)^2

OpenStudy (anonymous):

kk thanx!(: